OpenStudy (anonymous):

write sin3x-sin4x as a product

OpenStudy (anonymous):

\[\sin(a) - \sin(b) = 2\cos(\frac{a+b}{2})\sin(\frac{a-b}{2})\] Do the rest by yourself :)